# Estimating Knowledge in Large Language Models Without Generating a Single Token

### v1

@@ 5.3 Error Analysis
In §B, we include results for the other models, which follow the same trends in Figure 4 and Figure 5. We also provide the scatter plots for probes trained on the different KEEN features and baselines, all demonstrating the same linear relations.
@@ 5.4 Feature Analysis for VP-25 and VP-50
Figure 6: Difference, per subject, in the median rank of tokens with negative weight and tokens with positive weight. Pythia 12B VP-25 and VP-50 show the trade-off between interpretability and performance – though the median ranks of negative weight tokens are lower on average than positive weight tokens in VP-50, there is still a clear split in both accuracy groups.
@@
We analyze the most influential features of the KEEN QA VP probes to understand which tokens contribute most to predicting average QA accuracy. Our goal in this analysis is to identify tokens that either increase or decrease predicted QA accuracy, and to determine whether they are promoted in the representations of subjects with high and low QA accuracy, respectively. As a concrete example, for subjects with low QA accuracy, we expect tokens that decrease QA accuracy to generally be ranked lower in the subject representations than tokens that increase QA accuracy. Since the input normalization scheme described in §3.1 normalizes a token’s logit in a given hidden state by its magnitude across the other subjects (not with respect to the other tokens in the hidden state), we can interpret the weight learned by the KEEN QA VP probe for each token as its direction and magnitude of influence on the predicted score.
First, we identify the tokens associated with the largest absolute weights in the KEEN QA VP probes, as they are most influential on the predicted score. Next, we compare the median rank of tokens with negative weights to those with positive weights in the vocabulary projections of subjects with high QA accuracy (1.0) and low QA accuracy (0.0). Figure 6 shows that for low QA accuracy subjects, the median rank of negative weight tokens is generally lower than that of positive weight tokens. Conversely, for high QA accuracy subjects, the median rank of negative weight tokens is generally higher than that of positive weight tokens. This opposing trend in the two accuracy groups indicates that there is a small set of tokens which hold signals for differentiating between subjects the model knows a lot about and those it knows less about.
We provide these important tokens in Table 5. Tokens assigned positive weight are related to meaningful concepts while tokens assigned negative weight are often numbers, abbreviations, or suffixes. A possible interpretation of this semantic difference between positive weight tokens and negative weight tokens is that the hidden states of low accuracy subjects encode less content, reflecting the model’s lack of knowledge about them.
@@ 7 Conclusion
While our approach successfully estimates the extent of the model’s knowledge about a subject, it does not identify specific gaps or clusters of knowledge. For instance, KEEN can estimate that the model will be 55% truthful when generating content about Napoleon, but it does not pinpoint that the model is unable to answer the specific question, What military academy did Napoleon attend?. An interesting direction for future work would be to develop a more fine-grained approach that predicts how knowledgeable the model is about specific aspects of the subject (e.g. military career of Napoleon) or identifies specific facts encoded in subject representations.
@@ Acknowledgements
We thank Amir Globerson for insightful discussions and feedback throughout the project. We also thank Ori Yoran and Gal Yona for providing useful feedback on this manuscript. This work was supported by the Tel Aviv University Center for AI and Data Science (TAD) and the Israeli Science Foundation.
@@ Appendix A KEEN Training Details
All KEEN QA and OEG probes were trained with the AdamW optimizer with weight decay 0.01, and batch size of 32.
KEEN QA hyper-parameters were optimized over the configuration combinations presented in Table 6.
Hyper-parameter Values
Learning Rate $10^{-3},5e^{-3},5e^{-4},10^{-4},5e^{-5}$
Epochs 100, 1K, 3K, 5K
Table 6: KEEN QA probe hyper-parameter configurations
KEEN OEG hyper-parameters were optimized over the configuration combinations presented in Table 7.
Hyper-parameter Values
Learning Rate $5e^{-4},10^{-4},5e^{-5}$, $10^{-5}$
Epochs 1K, 3K, 5K
Table 7: KEEN OEG probe hyperparameter configurations
The best hyper-parameters for QA and OEG KEEN probes are found in Table 8 and Table 9, respectively.
KEEN Hyper GPT2 Pythia Pythia LLaMA LLaMA Vicuna Vicuna
Probe Param XL 6B 12B 7B 13B 7B 13B
HS Epoch 3K 100 100 1K 1K 1K 100
LR $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-5}$ $10^{-4}$
VP Epoch 3K 100 3K 1K 1K 1K 3K
LR $10^{-5}$ $10^{-4}$ $10^{-5}$ $10^{-5}$ $10^{-5}$ $10^{-5}$ $10^{-5}$
VP-200 Epoch 3K 500 3K 3K 3K 3K 3K
LR $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
VP-100 Epoch 3K 500 3K 3K 3K 3K 3K
LR $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
VP-50 Epoch 3K 1K 3K 3K 3K 3K 3K
LR $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
VP-25 Epoch 5K 3K 3K 3K 3K 3K 3K
LR $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
VP-10 Epoch - - - 3K - - -
LR $10^{-5}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$ $10^{-4}$
FC Epoch - - - 1K - - -
LR - - - $10^{-5}$
ATTN Epoch - - - 1K - - -
LR - - - $10^{-5}$
Table 8: Best hyper-parameters for KEEN QA probes and baselines.
@@
LR $10^{-5}$ $5e^{-5}$
Table 9: Best hyper-parameters for KEEN OEG probes and baselines.
@@
Appendix B Additional Results
B.1 Mean Standard Error (MSE) for KEEN
@@
Table 12: MSE for KEEN QA probes and FActScore.
@@
B.2 Interpretability-Performance Tradeoff for KEEN VP-k
@@
B.3 QA correlation plots

### v2

@@ 5.3 Error Analysis
In §C.5 and §C.6, we include results for the other models, which follow the same trends in Figure 4 and Figure 5. We also provide the scatter plots for probes trained on the different KEEN features and baselines, all demonstrating the same linear relations.
@@ 5.4 Feature Analysis for VP-25 and VP-50
Figure 6: Difference, per subject, in the median rank of tokens with negative weight and tokens with positive weight. Pythia 12B VP-25 and VP-50 show the trade-off between interpretability and performance – though the median ranks of negative weight tokens are higher on average than positive weight tokens in VP-50, there is still a clear split in both accuracy groups.
@@
Identifying prominent features
We analyze the most influential features of the KEEN QA VP probes to understand which tokens contribute most to predicting average QA accuracy. Our goal in this analysis is to identify tokens that either increase or decrease predicted QA accuracy, and to determine whether they are promoted in the representations of subjects with high and low QA accuracy, respectively. As a concrete example, for subjects with low QA accuracy, we expect tokens that decrease QA accuracy to generally be ranked higher in the subject representations than tokens that increase QA accuracy (the highest rank is 0 which corresponds to the token with highest logit value). Since the input normalization scheme described in §3.1 normalizes a token’s logit in a given hidden state by its magnitude across all subjects (not with respect to the other tokens in the hidden state), we can interpret the weight learned by the KEEN QA VP probe for each token as its direction and magnitude of influence on the predicted score.
First, we identify the tokens associated with the largest absolute weights in the KEEN QA VP probes, as they are most influential on the predicted score. Next, we compare the median rank of tokens with negative weights to those with positive weights in the vocabulary projections of subjects with high QA accuracy (1.0) and low QA accuracy (0.0). Figure 6 shows that for low QA accuracy subjects, the median rank of negative weight tokens is generally higher than that of positive weight tokens. Conversely, for high QA accuracy subjects, the median rank of negative weight tokens is generally lower than that of positive weight tokens. This opposing trend in the two accuracy groups indicates that there is a small set of tokens which hold signals for differentiating between subjects the model knows more about and those it knows less about. We provide these important tokens in Table 5.
Mapping knowledge clusters and “holes”
Tokens assigned positive weight are related to meaningful concepts while tokens assigned negative weight are often numbers, abbreviations, or suffixes. A possible interpretation of this difference is that the hidden states of low accuracy subjects reflect “holes” in the model’s knowledge, thereby encoding less information and promoting less semantically meaningful tokens. These negative weight tokens may help identify additional “holes” if similarly promoted in other subject representations. Conversely, positive weight tokens can be used to identify clusters of knowledge. Table 6 shows that Pythia 12B and Vicuna 13B encode more information about political figures, athletic players, or movies. This is evidenced by the $0.20-0.55$ higher than average QA accuracy for a sample of subjects with high logit values ($\geq 0.65$) for tokens like Senator, Player, Movie, and athlet.
Token Subject Logit QA
Pythia 12B Senator Rand Paul 0.84 0.55
Amin Amidu Sulemana 0.79 0.67
Beau Biden 0.67 0.54
Chelsea Clinton 0.68 0.58
Mean Entity 0.41 0.35
Players Charlton Athletic F.C. 0.84 0.60
Aston Villa F.C. 0.81 0.60
AC Milan 0.79 0.75
Moneyball 0.76 0.67
Mean Entity 0.44 0.35
Vicuna 13B Movie Attarintiki Daredi 0.86 0.71
Iron Man 3 0.83 0.63
Friday the 13th Part 2 0.76 0.88
National Treasure: Book of Secrets 0.76 0.86
Mean Entity 0.31 0.45
athlet Charlotte Hornets 0.97 1.00
Chariots of Fire 0.84 0.63
Olympique Lyonnais 0.74 0.60
The Cowboys 0.67 0.71
Mean Entity 0.41 0.45
Table 6: Examples of subjects from the full validation set with high logit values for influential positive tokens. Their high QA accuracies, relative to the Mean Entity, indicate knowledge clustered around the specific concepts: political figures, athletes/players, and movies.
@@ 7 Conclusion
While our approach successfully estimates the extent of the model’s knowledge about a subject, it does not identify the presence or lack of knowledge about specific facts. For instance, KEEN can estimate that the model will be 55% truthful when generating content about Napoleon, but it does not pinpoint that the model is unable to answer the specific question, What military academy did Napoleon attend?. An interesting direction for future work would be to develop a more fine-grained approach that predicts how knowledgeable the model is about specific aspects of the subject (e.g. military career of Napoleon) or identifies specific facts encoded in subject representations.
@@ Acknowledgements
We thank Amir Globerson for insightful discussions and feedback throughout the project. We also thank Ori Yoran and Gal Yona for providing useful feedback on this manuscript. This work was supported by the Tel Aviv University Center for AI and Data Science (TAD), the Israeli Science Foundation, and the Deutsch Foundation.
@@ A.1 Hyper-parameter Tuning and Resources
All KEEN QA and OEG probes were trained with the AdamW optimizer with weight decay 0.01, and batch size of 32, and learning rates were optimized over $1e^{-3},5e^{-3},5e^{-4},1e^{-4},1e^{-5},5e^{-5}$. The best hyper-parameters for QA and OEG KEEN probes are found in Table 7 and Table 8, respectively. The epochs represents the maximum number of epochs set for training since we did not implement early stop, however all evaluations are done on the checkpoint with the highest Pearson correlation on the validation set.
KEEN Hyper GPT2 Pythia Pythia LLaMA LLaMA Vicuna
Probe Param XL 6B 12B 7B 13B 13B
HS Epoch 3K 100 100 1K 1K 100
LR $1e^{-5}$ $1e^{-4}$ $1e^{-4}$ $1e^{-5}$ $1e^{-5}$ $1e^{-4}$
VP Epoch 3K 100 3K 1K 1K 3K
LR $1e^{-5}$ $1e^{-4}$ $1e^{-5}$ $1e^{-5}$ $1e^{-5}$ $1e^{-5}$
VP-200 Epoch 3K 500 3K 3K 3K 3K
LR $1e^{-5}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
VP-100 Epoch 3K 500 3K 3K 3K 3K
LR $1e^{-5}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
VP-50 Epoch 3K 1K 3K 3K 3K 3K
LR $1e^{-5}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
VP-25 Epoch 5K 3K 3K 3K 3K 3K
LR $1e^{-5}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
VP-10 Epoch - - - 3K - -
LR $1e^{-5}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$ $1e^{-4}$
FC Epoch - - - 1K - -
LR - - - $1e^{-5}$
ATTN Epoch - - - 1K - -
LR - - - $1e^{-5}$
Table 7: Hyper-parameters for KEEN QA probes and baselines.
@@
LR $1e^{-5}$ $5e^{-5}$
Table 8: Hyper-parameters for KEEN OEG probes and baselines.
@@
A.2 Choice of Input Layer Configuration
Our choice of choosing representations from the upper middle layers and the last subject position were motivated by previous work (Geva et al., 2023; Meng et al., 2022; Hernandez et al., 2024a). The focus on the last subject position is because in auto-regressive models, earlier positions cannot capture the full subject name. To systematically defend this decision, we repeated experiments recorded in Table 4 with LLaMA2 7B and Pythia 12B and trained new KEEN QA VP probes using the following layers as features: first 3 layers (Early) , last 3 layers (Late), 1 layer upper-intermediate layer (One) (22, 25), 5 upper-intermediate layers (Five) ([19, 24), [22, 27)). Correlation values between KEEN estimates and QA accuracy are statistically significant ($p\leq 2e^{-50}$), and provided in Table 10.
Model Early Late One Five KEEN VP
Pythia 12B 0.50 0.64 0.60 0.64 0.64
LLaMA2 7B 0.45 0.63 0.64 0.64 0.64
Table 10: Pearson correlation values of QA probes trained on hidden representations from various layer configurations.
We find a clear benefit to using upper-intermediate to late layers over early layers and no improvement when averaging over multiple upper intermediate layers. Using the late layers has comparable correlation values as using the upper intermediate layers, which supports previous findings that knowledge is aggregated in the subject’s hidden states of the upper intermediate layers (Geva et al., 2023; Meng et al., 2022).
A.3 Discussion of KEEN Efficiency and Training Time Statistics
KEEN has two key efficiency advantages. (1) Though KEEN probes are only trained on a small set of entities, they generalize well, which is especially significant considering that models are trained on millions of entities. Estimating knowledge with KEEN, for subjects not used to train the probes, requires a single inference pass and the execution of the KEEN probe. In contrast, the computational cost of query-based evaluation requires multiple inference passes per query. With every query requiring 1-5 passes (due to tokenization) and an average number of 5.3 queries per subject, query-based evaluation takes 5.3-27 times longer. (2) KEEN can be applied ubiquitously to all subjects since it does not require labeled data, beyond a small sample for training. Labeling is a major hurdle, especially for OEG where complex processing and verification of the output is needed (Min et al., 2023). Query-based evaluation approaches also require access to knowledge bases (KB) which inherently restricts the subjects that can be evaluated to those in the KB.
To demonstrate the training time benefits of KEEN, in Table 11 we report the epoch and practical length of training for KEEN QA probes used in our evaluations.
Table 11: Practical number of training epochs and times of KEEN QA probes.
Model Probe Best Training
Epoch (Min)
GPT2 XL VP 110 5
HS 3K 20
Pythia 12B VP 279 30
HS 100 1
Pythia 6B VP 44 2.5
HS 100 8
LLaMA2 7B VP 167 5
HS 1K 10
LLaMA2 13B VP 1K 1
HS 321 16.5
Vicuna 13B VP 100 1
HS 202 5
Appendix B Dataset Details
B.1 Question Answering Dataset
Examples are sampled as (subject, predicate, object) triplets from Wikidata, and questions are formed using manually written templates with placeholders for the subject and answer entity type (from the “instance of” property in Wikidata). We form directed questions using answer entity types to decrease the ambiguity of the expected answer i.e. instead of “Where was Barack Obama born?” we ask “In what city was Barack Obama born?”. Also, to increase the coverage of knowledge per entity, we extended the PopQA dataset to include an additional 26 relations to a total of 42 relations. Examples of answer templates and questions are found in Table 12. The train split consists of 2,223 entities and 12,324 questions, validation split has 555 entities and 3,003 questions, and test split has 694 entities and 3,821 questions.
Relation Question Template Entity Count
genre What genre is [subj]? 1979
country of origin What is the country of origin of [subj]? 1574
director Who was the director of [subj]? 1196
screenwriter Who was the screenwriter of [subj]? 1174
producer Who was the producer of [subj]? 1163
occupation What is [subj]’s occupation? 1092
color What color is [subj]? 1044
composer Who was the composer of [subj]? 1041
place of birth In what [obj_type] was [subj] born? 977
country of citizenship What is [subj]’s country of citizenship? 922
country In what country is [subj]? 913
languages spoken, written or signed What language does [subj] speak? 632
sport What sport does [subj] play? 503
language of work or name What is the language of [subj]? 493
capital What is the capital of [subj]? 482
author Who is the author of [subj]? 452
performer Who is the performer of [subj]? 361
educated at What is the alma mater of [subj]? 359
place of death In what [obj_type] was [subj] born? 343
followed by What [obj_type] follows [subj]? 332
father Who is the father of [subj]? 327
religion or worldview What is the religion of [subj]? 276
member of sports team What sports team does [subj] play for? 270
record label What is the record label of [subj]? 267
mother Who is the mother of [subj]? 250
position played on team / speciality What sports position does [subj] play? 241
spouse Who is the spouse of [subj]? 218
participant in In what sports event did [subj] participate in? 218
publisher Who is the publisher of [subj]? 217
sibling Who is the sibling of [subj]? 212
child Who is the child of [subj]? 211
capital of What is [subj] the capital of? 211
native language What is the native language of [subj]? 172
religion or worldview What is the religion of [subj]? 168
member of political party What is the political party associated with [subj]? 135
work location In what [obj_type] does [subj] work in? 110
country for sport What country does [subj] play for? 92
headquarters location In what [obj_type] are the headquarters of [subj] located? 80
league What sports league does [subj] play in? 74
lyrics by Who wrote the lyrics of [subj]? 70
consecrator Who is the consecrator of [subj]? 33
editor Who is the editor of [subj]? 12
Table 12: Templates used for generating questions for the QA dataset with the counts of attributed entities.
B.2 Open-Ended Generation Dataset
The unlabeled 500 subject FActScore dataset was used to evaluate the KEEN OEG probes, and the transfer of KEEN QA probes to the OEG setting.
B.3 Model Hedging Behavior
The experiment described in §5.1 assessed the correlation between the fraction of queries for which the model exhibited hedging behavior and the KEEN score for a given entity. To determine the hedging fraction, the model was prompted with a set of common question-answer pairs about the entity, and the proportion of responses containing an exact match with some hedging phrase was calculated. Hedging phrases were identified through manual analysis of model responses and included the expressions ‘‘nobody knows,’’ ‘‘I’m sorry,’’ ‘‘I can’t seem to find the answer,’’ ‘‘could you help me,’’ ‘‘can anyone help me,’’ ‘‘I’m not sure,’’ ‘‘I don’t know,’’ ‘‘I’m not entirely sure,’’ ‘‘could you please provide more,’’ ‘‘could you provide more information,’’ ‘‘provide more context,’’ and ‘‘clarify your question.’’
Appendix C Additional Results
C.1 Mean Standard Error (MSE) for KEEN
@@
Table 14: MSE for KEEN QA probes and FActScore. The MSE scores for FC are omitted because we do not compute KEEN probes on this feature.
@@
C.2 Interpretability-Performance Tradeoff for KEEN VP-k
@@
C.3 Recovering QA accuracy of non-target subjects through patching
The source model used in both patching experiments was LLaMA2 7B, fine-tuned for 100 epochs on passages from the Wikipedia page about Adil Shamoo. For both experiments, we analyzed 85 non-target subjects with $\geq 5$ questions whose QA accuracy decreased by at least 5% and KEEN scores either increased or stayed constant post fine-tuning. Representations from upper intermediate layers (20-23) of the source model were patched into the penultimate layer (30) of the target model. In the PT layer experiment, the target model was pre-trained LLaMA2 7B, while in the FT subj. experiment, the source and target models were both the fine-tuned LLaMA2 7B model mentioned above. In computing patched accuracy, a question was marked as correct if patching from any source layer recovered the correct answer. Formally, patched QA accuracy per subject is computed according to $y_{QA,\text{patched}}^{(s)}:=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\left[\exists\ell\in\{20,21,22,23\}:\hat{a}_{i}^{\text{patched}(\ell)}=a_{i}\right]$, where $\hat{a}_{i}^{\text{patched}(\ell)}$ denotes the answer generated after patching from source layer $\ell$ into the target model, as described above.
While §5.2 discusses patching results aggregated across all 85 non-target entities, Table 16 presents per-entity results from both patching experiments, showing that QA accuracy can be effectively recovered from intermediate fine-tuned representations.
KEEN PT KEEN FT Acc. PT Acc. FT PT layer FT subj.
John, King of England 0.51 0.51 0.67 0.30 0.50 0.50
Maurice Le Boucher 0.32 0.14 0.29 0.14 0.43 0.29
The Deep 0.51 0.52 0.71 0.00 0.71 0.71
A Whole New World 0.45 0.50 0.40 0.20 0.40 0.80
Alexander the Great 0.56 0.59 0.60 0.36 0.55 0.55
Jaakko Laakso 0.42 0.43 0.63 0.13 0.50 0.38
Reds 0.48 0.49 0.86 0.29 0.71 0.71
WarGames 0.59 0.63 0.50 0.25 0.50 0.50
Charles V 0.33 0.34 0.54 0.23 0.77 0.62
Recovery 0.25 0.29 0.17 0.00 0.67 0.50
James I of Scotland 0.47 0.47 0.45 0.18 0.55 0.46
Mai Van Hoa 0.33 0.36 0.29 0.14 0.50 0.17
The Natural 0.48 0.50 0.63 0.13 0.50 0.50
The Bourne Legacy 0.67 0.69 0.67 0.22 0.67 0.56
Beck 0.48 0.51 0.44 0.22 0.78 0.56
Radek Opršal 0.46 0.48 0.57 0.14 0.43 0.29
Elena Romagnolo 0.42 0.46 0.50 0.17 0.33 0.33
Jaroslav Kocián 0.47 0.56 0.38 0.13 0.50 0.25
Table 16: Per-entity results from the patching experiments discussed in §5.2. The recovery of QA accuracy is evident from both the observed increase in accuracy over Acc. FT and similarity to Acc. PT after patching (PT layer, FT subj.). Also, the alignment of KEEN PT and FT scores with both Acc. PT and patched accuracy indicates that KEEN reflects the amount of information encoded in the representations rather than in the observed output of the fine-tuned model.
C.4 Statistical significance of correlations
Table 17, Table 18, Table 19 represent the p-values associated with the Pearson correlation values found in Table 2, Table 3, Table 4, respectively.
Model Freq. FC ATTN VP-50 VP HS
GPT2 XL ${2.54e^{-17}}$ ${8.24e^{-47}}$ ${7.70e^{-56}}$ ${1.72e^{-59}}$ ${9.80e^{-79}}$ ${1.73e^{-75}}$
Pythia 6B ${3.26e^{-19}}$ ${6.41e^{-75}}$ ${2.74e^{-85}}$ ${7.28e^{-90}}$ ${1.33e^{-106}}$ ${3.16e^{-106}}$
Pythia 12B ${1.03e^{-15}}$ ${4.46e^{-62}}$ ${2.25e^{-76}}$ ${2.53e^{-74}}$ ${8.72e^{-91}}$ ${6.46e^{-91}}$
LLaMA2 7B ${9.96e^{-15}}$ ${4.54e^{-49}}$ ${1.17e^{-70}}$ ${3.07e^{-56}}$ ${4.50e^{-89}}$ ${2.52e^{-89}}$
LLaMA2 13B ${7.81e^{-13}}$ ${6.77e^{-49}}$ ${1.50e^{-50}}$ ${1.98e^{-45}}$ ${3.43e^{-70}}$ ${1.74e^{-71}}$
Vicuna 13B ${5.15e^{-13}}$ ${1.29e^{-48}}$ ${1.68e^{-53}}$ ${3.00e^{-50}}$ ${2.07e^{-77}}$ ${3.25e^{-76}}$
Table 17: p-values for different models across KEEN QA probes.
Model Freq. FC ATTN VP-50 VP HS
Pythia 12B ${1.00e^{-2}}$ ${2.68e^{-6}}$ ${7.20e^{-11}}$ ${4.57e^{-9}}$ ${3.04e^{-10}}$ ${4.21e^{-11}}$
Vicuna 13B $2.05e^{-2}$ ${1.28e^{-3}}$ ${6.16e^{-6}}$ ${2.50e^{-4}}$ ${3.43e^{-6}}$ ${4.02e^{-6}}$
Table 18: p-values for different models across KEEN OEG probes
Model Freq. FC ATTN VP-50 VP HS
Pythia 12B ${2.74e^{-29}}$ ${1.00e^{-6}}$ ${8.43e^{-7}}$ ${1.76e^{-2}}$ ${4.92e^{-5}}$ ${3.25e^{-7}}$
Vicuna 13B ${2.28e^{-16}}$ ${4.80e^{-4}}$ ${1.39e^{-4}}$ ${7.84e^{-5}}$ ${9.18e^{-6}}$ ${2.12e^{-5}}$
Table 19: p-values for the correlation of KEEN QA probe and FActScore scores.
C.5 QA correlation plots